OpenStudy (anonymous):

Solve for x: −4|x + 5| = −16

OpenStudy (anonymous):

WILL GIVE MEDALS

OpenStudy (anonymous):

negative sign will cancelled

OpenStudy (anonymous):

then what

OpenStudy (anonymous):

|x+5|=16/4=4

OpenStudy (anonymous):

so whats the answer

OpenStudy (anonymous):

remove mode will lead to 2 solutions

OpenStudy (anonymous):

x+5=4 and x+5=-4 x=4-5=-1 and x=-4-5=-9

OpenStudy (anonymous):

x=-1 or x=-9

OpenStudy (anonymous):

these are my ooptions x = 21 over 4, x = −11 over 4 x = −1, x = 9 x = −1, x = −9 No solution

OpenStudy (anonymous):

thanks @g17

OpenStudy (anonymous):

hope u got it :P